40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Nov 4 23:55:24 2007, rana, Update, IOO, Friday's In-Vac work 
    Reply  Mon Nov 5 07:29:35 2007, rana, Update, IOO, Friday's In-Vac work 
Message ID: 61     Entry time: Sun Nov 4 23:55:24 2007     Reply to this: 62
Author: rana 
Type: Update 
Category: IOO 
Subject: Friday's In-Vac work 
On Friday morning when closing up we noticed that we could not get the MC to flash any modes.
We tracked this down to a misalignment of MC3. Rob went in and noticed that the stops were
still touching. Even after backing those off the beam from MC3 was hitting the east edge of
the MC tube within 12" of MC3.

This implied a misalignment of MC of ~5 mrad which is quite
large. At the end our best guess is that either I didn't put the indicator blocks in the
right place or that the MC3 tower was not slid all the way back into place. Since there
is such a strong stickiness between the table and the base of the tower its easy to
imagine the tower was misplaced.

So we looked at the beam on MC2 and twisted the MC3 tower. This got the beam back onto the
MC2 cage and required ~1/3 if the MC3 bias range to get the beam onto the center. We used
a good technique of finding that accurately: put an IR card in front of MC2 and then look
in from the south viewport of the MC2 chamber to eyeball the spot relative to the OSEMs.

Hitting MC2 in the middle instantly got us multiple round trips of the beam so we decided
to close up. First thing Monday we will put on the MC1/MC3 access connector and then
pump down.


Its possible that the MC length has changed by ~1-2 mm. So we should remeasure the length
and see if we need to reset frequencies and rephase stuff.
ELOG V3.1.3-